During the first 4 hours of a trip, you average 40MPH. How long must you travel at 60MPH to finish with an overall average speed of 55MPH?
Answer by Paul(988)

let x be the time
EQUATION:
40(4)+60x=55(x+4)
160+60x=55x+220
5x=60
x=12
HEnce, you must travel for 12 hours.
Paul.